OpenStudy (anonymous):

The surface area of a sphere is found using the formula SA=4(3.14)^2 , where r is the radius. If the surface area of a sphere is 64(3.14) square feet, what is the radius, in feet? A. –4 B. 16 C. 4 D. –16

OpenStudy (anonymous):

I know this, I did a portfolio on surface area. Give me a minute and I will look for my homework to help you, there's a little trick you can use.

OpenStudy (anonymous):

I have it for Prism, Pyramid, Cylinder and Cone. I will go back more to find area of a sphere.

OpenStudy (anonymous):

Prism: SA = (perimeter of base x height) + (Area of base) L+L+W+WxH 2xLxW Volume = LxWxH Volume of a Pyramid = Do the same thing you did for the volume of a prism only x 1/3 or divide by 3. Cylinder = SA = (2 x 3.14 x R x H) + (2 x 3.14 x R x R) Volume = 3.14x R*2xH Volume of a Cone = (Volume of a cylinder) x 1/3 or divide by 3. I will work on your answer in a moment. I have to find it in my book.

OpenStudy (anonymous):

I got 4. So C.